Grade 11General Physics

A vector of magnitude 30 and direction eastward is added with another vector of magnitude 40 and direction northwards find the direction of resultant with east.

Shaunak Deshpande
8 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Resultant vector = 30i + 40j
Magnitude = √(30² + 40²)
= 50
Direction
Tanθ = 40/30 = 4/3
θ = 37° with East
 
Regards
Arun (askIITians forum expert)